Prove that a sequence in a metric space cannot converge more than one limit


1
Expert's answer
2021-08-03T07:07:49-0400

Let us prove that a sequence "(x_n)_{n\in\N}" in a metric space "(X,d)" cannot converge more than one limit using the method by contradiction. Suppose that the sequence "(x_n)_{n\in\N}" converges to "x" and also it converges to "y\ne x." Let "\rho=d(x,y)." Since "x\ne y," we conclude that "\rho>0." Let "\varepsilon=\frac{\rho}{2}." By definition of limit there exists "n_x\in\N" such that "d(x_n,x)<\varepsilon" for all "n\ge n_x." Also there exists "n_y\in\N" such that "d(x_n, y)<\varepsilon" for all "n\ge n_y." Let "n_0=\max\{n_x,n_y\}." Then "d(x,y)<d(x,x_n)+d(x_n,y)<\varepsilon+\varepsilon=2\varepsilon=\rho" for all "n\ge n_0," a contradiction with "d(x,y)=\rho." We conclude that a sequence in a metric space cannot converge more than one limit.